solve using Newton Method
Numerical TJ Find the mulliplicity of rock 1 f(x) = (x-1)2 inx 2 Find the order of convergence Pit en 9 digits accuracy 13 give the roo
Answer #1

O hiven: f() = (x-1)2 Inco) -) f(1) = 0 f(x) = 2(x-1) In (30+ (x-1) 2 =) f(1) = 0. t(x) = 2 ln(x) + 2(x-1) X (2(x-1))(x-1)3